### v2

# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

Abstract
Human Activity Recognition (HAR) is a longstanding problem in AI with applications in a broad range of areas, including healthcare, sports and fitness, security, and more. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundation models, can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g., in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. To alleviate the need for labeled data, we introduce our comprehensive Fitness Multimodal Activity Dataset (FiMAD) in this work, which can be used with the proposed pre-training method MuJo (Multimodal Joint Feature Space Learning) to enhance HAR performance across various modalities. FiMAD was created using YouTube fitness videos and contains parallel video, language, pose, and simulated IMU sensor data. MuJo utilizes this dataset to learn a joint feature space for these modalities. We show that classifiers pre-trained on FiMAD can increase the performance on real HAR datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH. For instance, on MM-Fit, we achieve an Macro $F_{1}$-Score of up to 0.855 when fine-tuning on only 2% of the training data and 0.942 when utilizing the full training set for classification tasks. We have compared our approach to other self-supervised ones and showed that, unlike them, ours can consistently improve on the baseline network performance as well as provide a better data-efficiency.
{NoHyper}00footnotetext: Equal contribution.
1 Introduction
Figure 1: The pipeline depicts the construction process of FiMAD and the training of MuJo for multimodal joint feature space learning. The asterisk (*) indicates that the input is being pre-calculated (frozen) and not optimized during the training process.
Perceiving and interpreting human activities is a core functionality of intelligent systems with applications in a broad range of areas, including healthcare Tan et al. (2021), sports and fitness Host and Ivašić-Kos (2022); Nadeem et al. (2020), and security Sunil et al. (2021), among others. A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in Computer Vision, particularly with Large Vision-Language Models (LVLMs) Radford et al. (2021); Achiam et al. (2023), have enabled impressive performance when given sufficiently high-quality images. In many cases, merely feeding images to publicly available tools such as GPT-4 can result in accurate description of scenes involving complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis. Unfortunately, achieving acceptable recognition performance with such sensors is a much more challenging problem due to two factors, the first one being information ambiguity. An IMU (a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture. This problem is heightened by the second factor, lack of training data. The success of LVLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations. Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al. (2024), labeled training data remains scarce.
The work presented in this paper, which aims to address these issues, is founded on three key observations:
  1. 1.
    While the input modalities availability during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
  2. 2.
    Representation learning facilitates knowledge transfer between modalities by aligning embedding spaces, in particular when synchronized training data is available.
  3. 3.
    Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube Kwon et al. (2021) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods like OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019) are used to generate skeletal data of human activities. This enables the creation of synchronized video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a Multimodal Joint Feature Space Learning method (MuJo) that learns a joint representation between videos, video-derived poses, virtual IMU data generated from videos, and textual descriptions by employing pair-wise contrastive learning, as depicted in Figure 1. MuJo aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. For that, we provide extensive evaluation on four publicly available HAR datasets, namely MM-Fit Strömbäck et al. (2020), MyoGym Koskimäki et al. (2017), MotionSense Malekzadeh et al. (2018), and MHEALTH Banos et al. (2015). For instance, on the MM-Fit dataset in the multimodal case (encompassing all of the aforementioned modalities as input), our method improves the recognition rate (Macro $F_{1}$-Score) from 0.897 to 0.942 when using all available training data. Even when utilizing only 2% of the training data, it enhances the recognition rate from 0.790 to 0.855.
Overall, the key contributions of this work can be summarized as follows:
  1. 1.
    Development of the Fitness Multimodal Activity Dataset (FiMAD) [^1] Available after publication at link., a comprehensive dataset containing parallel video, language (including video clip labels, instructor utterances, and detailed exercise descriptions generated by GPT-3.5 Turbo), pose data, and simulated IMU data points.
  2. 2.
    Pe-training of MuJo, a multimodal joint space model, using the FiMAD dataset. MuJo integrates four different modalities: text, video, pose, and virtual sensor data, leveraging contrastive learning inspired by CLIP Radford et al. (2021).
  3. 3.
    Evaluation of the utility and generalizability of MuJo for real-world accelerometer data on the MM-Fit, MyoGym, MotionSense, and MHEALTH datasets. On the MM-Fit dataset, we additionally conducted experiments with pose, video, and multimodal inputs. The benefits of pre-training MuJo on FiMAD, especially when limited data is available for the target task, are analyzed. Furthermore, we analyze the zero-shot performance of our method and compare unimodal versus multimodal training on all four datasets.
2 Related Work
Unimodal and multimodal activity recognition.
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task, and they can be broadly categorized into two groups: hand-crafted feature-based and deep learning-based methods. Hand-crafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition Kwapisz et al. (2011), features such as mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition Strömbäck et al. (2020), being combined with RNNs in some cases Ordóñez and Roggen (2016). Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
Multimodal activity recognition, on the other hand, recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al. (2022). These features can then be combined using fusion methods such as early fusion or late fusion Duhme et al. (2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment Plananamente et al. (2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems Islam et al. (2022).
Representation learning with proxy tasks.
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion Seo et al. (2021); Sun et al. (2019), matching/alignment Li et al. (2020); Miech et al. (2020), and ordering tasks Li et al. (2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods, such as concatenating learned uni-modal feature vectors, have been extensively explored to enhance the performance of downstream tasks Hu et al. (2018); Strömbäck et al. (2020). These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. More advanced methods have also been proposed, including a tensor fusion network Zadeh et al. (2017) and a transformer-based co-attention mechanism for efficient cross-modality learning Lu et al. (2019).
Proxy task were also developed for the field of Human Activity Recognition with weareable devices including IMUs, where unlabelled IMU data is used to pre-train better representations. Those approaches include Multitask Saeed et al. (2019), Auto-encoder Haresamudram et al. (2019) and SimClr Tang et al. (2021). Multi-task Saeed et al. (2019) proposes auxiliary tasks of binary classification for extracting useful features for the downstream task. A multi-task temporal convolutional neural network (CNN) is trained to classify whether the input signal was transformed or not for each of a set of transformations such as scaling, rotating the signal, etc. The authors utilize a large unlabelled accelerometer dataset for self-supervised learning and transfer learned knowledge to a small, labelled activity recognition dataset. Auto-encoder Haresamudram et al. (2019) is the application of auto-encoders where an encoder network reduces the dimensionality of the data while a decoder one brings the representation back to its original dimensions, with the whole system learning to reconstruct the original signal. SimClr Tang et al. (2021) learns general representations from a dataset where a model is trained to match different views instances created from the same signal by contrasting them. The positive pair is obtained by the transformation of the input window, while the negative pair is formed by the rest input windows of the same batch. Then, the contrastive loss pulls the positive pair together while pulling the negative pair far away from each other in the representation space.
Representation via contrastive learning.
Contrastive learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar and dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP Radford et al. (2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning has been extended to other modalities as well. For example, video-text data has attracted interest in studies focused on the shared space of video-text features Xu et al. (2021); Xue et al. (2022) and text-to-video generation Singer et al. (2022). MotionCLIP Tevet et al. (2022) trained an encoder to find the appropriate embedding of an input sequence in CLIP space and a decoder to generate the most fitting motion for a given CLIP space. Following this, IMU2CLIP Moon et al. (2022) was proposed to learn a joint space that aligns IMU motion sensor recordings with video and text by projecting them into a shared representation space.
Studies in action recognition Li et al. (2020); Wang et al. (2021); Sun et al. (2022); Morshed et al. (2023) have also significantly enhanced the field of fitness activity recognition, despite traditionally focusing on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”. Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning, as exemplified by initiatives like HERO Li et al. (2020) and ActionCLIP Wang et al. (2021).
The achievements discussed in the aforementioned studies have shown that contrastive learning has a significant impact on extracting a joint space with similar and dissimilar pairs. Consequently, the contrastive learning methodology employed in our study aims to acquire knowledge of the feature space pre-training to fitness activities. Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al. (2023). Thus, our model and data provide a cross-modal classification opportunity.
3 Dataset Collection and Construction
Annotating video descriptions is time-consuming and expensive. Thus, we create our dataset from a comprehensive and varied collection of videos with temporally localized chapter information, including chapter titles and start times, specifically for fitness activities. YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Those chapters are manually added by the uploader or automatically generated. Each chapter is a continuous, non-overlapping segment that divides the video into a video clip and its corresponding textual description.
We manually collect 1388 fitness videos with pre-annotated video chapters. Additionally, we retrieve the captions automatically generated by YouTube. All our videos show a tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). The trainers are primarily performing HIIT home workouts, focusing on bodyweight exercises with no or less equipment. The exercise videos provide instructions for the activities while showing the activities clearly. The annotation does not include predefined categories or labels. Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Based on the temporal boundaries of the chapters within each video, we split the videos in shorter activity clips along with their corresponding caption segments. Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter annotations. Hence, we only keep exercises shorter than 2 minutes and discard longer exercises. After that, the dataset consists of 10,695 (video, chapter, caption)-samples.
We convert all 10,695 exercise videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation. To enhance the dataset with diverse modalities, we use IMUTube Kwon et al. (2021) to obtain pose and virtual sensor (accelerometer and gyroscope) features. As a result, our dataset includes (video, pose, sensor, text)-pairs for each input. For video and text we extracte multiple features. Video features include using HERO Li et al. (2020), which is utilizing either CLIP Radford et al. (2021) or S3D Xie et al. (2018) embeddings. For text, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list. Further, we use GPT-3.5 Turbo to generate detailed descriptions of the exercises. We then extract multiple text features with various models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT, as well as language features and motion tokens extracted with the Text2Motion Transformer Guo et al. (2022). We apply IMUTube to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3D pose lifting. Since IMUTube fails to extract sensor data for 884 exercise videos, we exclude these sequences. We use the remaining 9811 exercises as our dataset and split it into 8,227 train and 1,584 validation activity clips. The split is based on the level of the 1,388 original fitness videos, to ensure video independence. We do not create a test set because the dataset is used solely for pre-training, not for testing. Since workout session videos typically start and end with explanations given by the trainer, we discard the initial and final 15% [^2] This value has proven to be effective in experimental investigations. of each activity clip. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances are generated using sliding windows $x_{i}$ of 100 frames in length (i.e., 2 seconds) with 50 frames overlap from the videos. We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements. In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, virtual IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.
4 Method
Our method leverages multimodal information for HAR pre-training. In this Section, we outline the task and its formulation, along with detailing the multimodal representation methods employing CLIP Radford et al. (2021)-like contrastive learning. Our objective is to improve HAR by employing contrastive learning across multiple modalities. The fundamental idea is that the extracted information from different modalities for a short video interval should be similar, as they represent the same activity from various perspectives. Each instance is a window, and for each one, we extract features for every modality. This means our data consists of $\mathcal{X}^{m}=\{x^{1}_{m},...,x^{N}_{m}\}$, where $x^{i}_{m}$ is the feature for modality $m$ at window $i$. We use as modalities:
  • $\mathcal{X}^{pose}$: The sequence of 3D poses for the subject extracted with ImuTube Kwon et al. (2021) utilizing OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019).
  • $\mathcal{X}^{video}$: Video features extracted with HERO Li et al. (2020) utilizing CLIP Radford et al. (2021) embeddings.
  • $\mathcal{X}^{acc}$: Simulated accelerometer data generated by ImuTube.
  • $\mathcal{X}^{text}$: Embeddings extracted using OpenAI’s ADA model from YouTube’s chapter information with a lightweight normalization process removing a few special symbols and replacing abbreviations.
Each modality $m$ has its own encoder network $e_{m}$ and projection layers $p_{m}$. Thus, $rep_{m}(x_{m})=p_{m}(e_{m}(v_{m}))$ serves as our representation for a vector $v_{m}$ of modality $m$ in our shared representation space. We learn this shared space between modality representations using the InfoNCE loss. First, for all possible $(m_{a},m_{b})$ modality pairs, we compute similarities using
${sim(v_{a},v_{b})}=\frac{{rep_{m_{a}}(v_{a})}\times{rep_{m_{b}}(v_{b})}^{\top}}{\lVert{rep_{m_{a}}(v_{a})}\rVert\cdot\lVert{rep_{m_{b}}(v_{b})}\rVert}$ (1)
Then, those similarities are softmax-normalized within the batch to calculate the loss between the two modalities
$\mathcal{L}_{\text{c}}(x_{a},x_{b})=-\sum_{i}\log\frac{\text{exp}(sim(x^{i}_{a},x^{i}_{b})/\tau)}{\sum_{j}\text{exp}(sim(x^{i}_{a},x^{j}_{b})/\tau)},$ (2)
where $x_{a}$ and $x_{b}$ correspond to aligned batch data for modalities $a$ and $b$, respectively, and $\tau$ is the temperature hyper-parameter. Since we calculate losses between all modalities, our total loss amounts to
$\mathcal{L}_{total}=\frac{1}{\binom{M}{2}}\sum_{m_{a},m_{b}}(\mathcal{L}_{c}(x_{m_{a}},x_{m_{b}})+\mathcal{L}_{c}(x_{m_{b}},x_{m_{a}})),$ (3)
where $M$ is the total number of modalities (in our case 4). An example can be seen in Figure 1.
Regarding the structure of our encoders, we utilize a model with a single fully connected layer for $e_{text}$ and two fully connected layers for $e_{video}$. For $e_{pose}$ and $e_{sensor}$, we employ the CNN encoder architecture as described in Strömbäck et al. (2020), making slight adaptations to accommodate our input shapes. The projection $p_{m}$ is a two-layer fully connected model for all modalities $m$, mapping them to feature vectors of length 1280.
Hyperparameter and model optimization.
We applied a random search approach for hyperparameter selection. Our hyperparameters include learning rate, number and size of projection layers, dropout rates, and activation functions. Furthermore, we experimented with various configurations of the loss function, specifically comparing the effect of using the text modality as an anchor, i.e., only considering sensor-text, pose-text, and video-text loss pairs, versus incorporating all possible modality combinations into the loss. Encompassing all modality pairs together yielded superior results. We explored weighting the text modality higher than the others in the loss, but it did not have any observable effect on performance. We also explored the use of S3D-based video embeddings, the gyroscope as sensor input, and the other text modalities and features described in Section 3 as inputs. However, since those configurations did not yield optimal performance during the hyper-parameter search, we have omitted those results for the sake of brevity.
5 Evaluation Methodology
5.1 Evaluation Datasets
We conduct experiments on the MM-Fit Strömbäck et al. (2020), MyoGym Koskimäki et al. (2017), MotionSense Malekzadeh et al. (2018), and MHEALTH Banos et al. (2015) datasets. MM-Fit is designed for multimodal fitness activity recognition and combines data from multiple modalities, including video, pose estimations, and various sensors. The dataset includes 21 workout sessions conducted by ten subjects, each performing ten different fitness exercises. The MyoGym dataset comprises 30 distinct gym exercises focused on upper body training by ten individuals, each equipped with a sensor on their right forearm. The MotionSense dataset consists of sensor data collected from 24 participants carrying an iPhone 6s in their front pocket, covering six different activities (walking downstairs, walking upstairs, sitting, standing, and jogging). The MHEALTH dataset includes twelve physical activities (such as climbing stairs, cycling, and running) performed by ten subjects, with data captured by three sensors positioned on the chest, right wrist, and left ankle. All datasets, except for MotionSense, include a "No Activity" category (NULL class) in addition to the specified activities.
5.2 Unimodal Classification on Real-World Accelerometer Data
We evaluate the performance of our pre-trained representation on real-world accelerometer data using all four datasets. For MM-Fit, we use the accelerometer data from the smartwatches worn on the left and right wrists, which we had to downsample from 100 Hz to 50 Hz via linear interpolation to match our input frequency. For MyoGym, MotionSense, and MHEALTH, we use accelerometer data from all available sensors, which are already at 50 Hz, allowing us to employ the data directly without additional preprocessing. For all four datasets, we apply $Z$-score normalization and use a sliding window of 100 samples in length (i.e., 2 seconds) with 50 samples overlap to generate instances. Furthermore, since each target dataset has different sensor placements, we have to pre-train a separate MuJo model for each one. Specifically, during pre-training our models, we use the virtual sensor positions generated by IMUTube that best match the sensor positions of the respective target dataset. Each of the four models is pre-trained on our FiMAD dataset incorporating all four input modalities: text, video, poses, and virtual accelerometer data.
We conduct our experiments based on the standard experimental methodology in the HAR community, i.e., testing on unseen users and evaluating using the Macro $F_{1}$-Score, a metric that is well-suited for handling class imbalance. For MM-Fit, we use the training, validation, and cross-subject test sets as described in Strömbäck et al. (2020). Since no fixed splits are provided for MyoGym, MotionSense, and MHEALTH, we employ $k$-fold cross-validation with $k=5$. Additionally, we consider two scenarios: training the classifiers with and without the inclusion of the NULL class. This is relevant because the NULL class is not part of our pre-training dataset but is present in real-world scenarios and included in all evaluation datasets except for MotionSense.
For classification, we utilize a classifier architecture $C$ comprising our sensor encoder $e_{sensor}$, projection $p_{sensor}$, and a classification head $cl$ consisting of 2 layers. We contrast three scenarios: In the baseline setting, all weights are initialized randomly, meaning the model lacks any pre-trained knowledge. In the other two scenarios, the encoder and projection head are initialized with weights from our pre-trained model. In one case, the weights of the encoder and projection head are fixed (frozen), whereas in the other case, they are trainable and we fine-tune them during the classifier’s training process. We train with different proportions of the training data, ranging from one to 100 percent, with the training data being sampled randomly per class to ensure balanced representation. To guarantee comparability of all results, we train each classifier $C$ with the same settings: Training is conducted for a maximum of 200 epochs, incorporating early stopping after 25 epochs without improvement. We employ a weighted cross-entropy classification loss to address class imbalance, given by:
$\mathcal{L}_{CL}\;\;=\;\;-\sum_{i=1}^{n}w_{y_{i}}y_{i}\log(C(x_{i}))$ (4)
where $n$ is the number of classes, $y_{i}$ is a one-hot encoded ground truth activity label and $C(x_{i})$ is the predicted label vector of $C$ for input $x_{i}$, that is, the application of our $cl$ classification layers to obtain $cl(p(e(x_{i})))$. The weight $w_{y_{i}}=\frac{\max{(M)}}{m_{y_{i}}}$ for class $y_{i}$ is calculated by dividing the number of data samples of the largest class by the number of instances in class $y_{i}$ itself. To ensure validity of the results, we train 20 models independently for each configuration, with the training data being randomly selected anew for each run (for proportions less than 100%) and the non-pre-trained weights being reinitialized randomly.
5.3 Proxy task comparisons
We compare our approach with Multi-task Saeed et al. (2019), Auto-encoder Haresamudram et al. (2019) and SimClr Tang et al. (2021). All approaches were trained with the same simulated IMU data as our approach and use the same encoder architecture as ours, for fair comparison. Regarding other configurations SimClr used random rotations as augmentation uniformely sampled in $(-\pi,\pi)$ range. The representation is obtained from our projection layer through two dense layers of size 256, 128 and 128, respectively, with the first two being followed by a relu activation. It was trained with the NTXent loss using l2 norm. For Multitask we have selected as augmentations noising, scaling, negation, horizontally flipping, permutation and channel shuffling. Noising consists of adding a value sampled from a normal distribution with mean zero and standard deviation of 0.05 while scaling consists of multiplying by values drawn from this same distribution. Negation, horizontally flipping and channel shuffling are self evident. Permutation means randomly splitting the time series into four parts and changing their order. Every task has a head that takes as input the encoder’s output and consists of one fully connected layer with 256 size and relu activation, followed by another fully connected layer with also 256 size but sigmoid activation. For Auto-encoder the representation is passed to a convolutional decoder that re-creates the original signal. This decoder consists of first a linear layer to bring the projection to $(32,n\_sensors,\frac{win\_size}{4})$ where $n\_sensors$ is the number of sensors in the dataset and $win\_size$ is the original window size. This is followed by batch norm, relu and then three blocks of convolution transposed, $0.05$ dropout, relu and upsampling. A kernel size of $(1,11)$ was used and the channels change following the sequence $(32,16,8,orig)$, with $orig$ being the original number of sensor channels. To our knowledge, this is the first application of such methods to simulated sensor data, as well as their first use in a multi-sensor scenario. While all those methods could be added to our framework easily, we selected them as comparison to highlight how our method can learn from extra modalities only available at training time beyond what is present at the simulated sensor data itself (which is the only input for those methods).
5.4 Multimodal Classification on MM-Fit
We evaluate the multimodal capabilities of our system using the MM-Fit dataset. Since MM-Fit comes along with videos and poses at 30 Hz, we upsample both modalities to 50 Hz. We extract HERO CLIP features from the videos following the approach outlined in Section 3. We use the sliding window method described in Section 5.2 to generate instances for pose and video inputs. We also employ the same classifier architecture $C$ but select the encoders $e_{m}$ and projections $p_{m}$ depending on the input modality. Specifically, we train individual classifier models for the pose and video input modality, using their respective encoders $e_{m}$ and projections $p_{m}$. Additionally, we experiment with multimodal input by providing all three modalities (sensor, pose, and video) to the model and concatenating the outputs from the respective projections before feeding them into the classifier head $cl$. We keep the same training setup as outlined for unimodal training in Section 5.2, perform training runs with and without the NULL class, and contrast the same three training scenarios: baseline, pre-trained, and pre-trained with frozen encoder and projection weights. For the pre-trained scenarios, we reuse the MuJo model from Section 5.2 that was pre-trained for the MM-Fit target dataset (i.e., allowing left and right wrist sensor placements).
{NiceTabular}
|c|c|c|ccc|ccc|
Dataset With “NULL” Method Mean Macro $F_{1}$-Score
Train Data: $2\%$ Train Data: $100\%$
Baseline ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Baseline ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Pre-trained ([Uncaptioned image])
[…]
Auto-encoder .$065\pm.016$$253\pm.019$$372\pm.043$$435\pm.024$
SimClr .$038\pm.006$$088\pm.016$$081\pm.022$$447\pm.043$
Multitask .$241\pm.025$$278\pm.029$$425\pm.052$$465\pm.043$
[…]
2-9
[…]
Auto-encoder .$116\pm.001$$204\pm.002$$242\pm.028$$344\pm.020$
SimClr .$046\pm.005$$094\pm.015$$102\pm.018$$352\pm.038$
Multitask .$194\pm.022$$233\pm.002$$339\pm.047$$376\pm.040$
[…]
Auto-encoder .$527\pm.042$$538\pm.039$$785\pm.343$$826\pm.033$
[…]
Multitask .$534\pm.083$$626\pm.099$$758\pm.079$$758\pm.079$
[…]
2-9
[…]
Auto-encoder .$321\pm.155$$362\pm.026$$514\pm.030$$539\pm.040$
[…]
Multitask .$313\pm.032$$407\pm.048$$414\pm.050$$547\pm.056$
MM-Fit MuJo .$423\pm.045$$474\pm.028$$492\pm.027$$544\pm.025$$557\pm.005$$580\pm.014$
Auto-encoder .$190\pm.061$$394\pm.070$$511\pm.009$$556\pm.020$
SimClr .$135\pm.015$$337\pm.036$$315\pm.016$$522\pm.023$
[…]
2-9
MuJo .$257\pm.023$$237\pm.022$$312\pm.015$$389\pm.028$$368\pm.006$$408\pm.010$
Auto-encoder .$232\pm.187$$243\pm.037$$283\pm.009$$382\pm.011$
SimClr .$112\pm.009$$223\pm.020$$202\pm.011$$356\pm.019$
[…]
SimClr .$640\pm.385$$699\pm.324$$729\pm.031$$850\pm.048$
Multitask .$598\pm.377$$691\pm.445$$725\pm.040$$834\pm.055$
Table 1: Comparison of classification results on accelerometer data varying the training amount using different methods for pre-training encoders. We also investigate the effect of freezing encoders during fine-tuning.
Figure 2: Classification performance across various training data fractions. The comparison includes the baseline model and models with a pre-trained MuJo encoder and projection (both with frozen and trainable weights) on accelerometer data for all evaluated datasets.
Figure 3: Classification performance across various methods and training data fractions, where each boxplot represents the results of 20 runs. The comparison includes the baseline model and models with a pre-trained encoder and projection with trainable weights on accelerometer data for all evaluated datasets.
5.5 Zero-Shot
[…]
6 Results and Discussion
6.1 Unimodal Classification on Real-World Accelerometer Data
Table 5.4 shows that for our method the pre-trained models with trainable weights consistently outperform the baseline with respect to Macro $F_{1}$-Score on all dataset, regardless of whether 2% or 100% of the training data is used. These findings indicate that multimodal pre-training via contrastive learning on virtually generated accelerometer data can effectively improve HAR performance on real-world sensor data. Regarding comparison with other approaches, we can see how our method is more data-efficient, as we outperform other approaches in almost all cases when 2% of the labeled training data is available. When 100% of the training data is available every method is dominant in a different dataset, but ours consistently improves on the baseline unlike other approaches. Moreover, our method presents a more consistent performance when freezing the encoder, showing the quality of the learned representation.
In general including the NULL class generally results in lower scores, but our trainable models still achieve higher scores than the baseline, even if this class was not part of our pre-training. Only for the MHEALTH dataset, when utilizing 100% of the data, the performance difference becomes marginal, with scores of 0.839 versus 0.837 without the NULL class and 0.582 versus 0.577 with the NULL class. However, when labeled data is limited, we perform significantly better than the baseline. For example, when utilizing only 2% of the train data, even the pre-trained classifiers with frozen encoder and projection layers outperform the baseline in all cases except for the MM-Fit dataset when the NULL class is included.
The boxplots presented in Figure 2 compare the baseline to classifiers pre-trained with MuJo, both with frozen and trainable encoders and projections. For various training fractions ranging from 1% to 100%, we observe that the performance of our frozen representations highly depend on the target dataset and whether the NULL class is included during training. For instance, when the NULL class is excluded, the frozen models perform comparable to the pre-trained trainable models on MYOGYM and only slightly worse on MM-Fit. On MYOGYM with the NULL class, the frozen classifiers perform comparably to the baseline, whereas in all other scenarios, they generally outperform the baseline when less training data is utilized, but perform worse when more data is included in the training. On the other hand, the pre-trained models with trainable weights constantly achieve higher scores than the baseline for all training fractions on all datasets, which demonstrates the effectiveness of our pre-trained representations for accelerometer-based activity recognition, and robustness to varying sensor placements and target domains.
We have also compared MuJo to the other pre-trained methods with different training data fractions, all with trainable encoders, see Figure 3. Here it is clear that our method is clearly the best in the MYOGYM dataset. Moreover, it outperforms all others up to 10% training fraction in all dataset other than MM-Fit, where only Multitask is better.
6.2 Multimodal Classification on MM-Fit
Table 6.2 presents the results of our experiments conducted for pose, video, and multimodal (concatenation of sensor, pose, and video) inputs, using both 2% and 100% of the training data, and with and without the inclusion of the NULL class. While the classifiers with our frozen pre-trained MuJo encoders and projections often perform worse than the baseline, our trainable models outperform the baseline in all cases except for the pose input without the NULL class when using the full training data. In general, for the pose input, the performance difference between these two training scenarios is only marginal, as evident from Figure 4. This discrepancy may be attributed to differences between the poses generated by IMUTube and that extracted by MM-Fit. Although both IMUTube and MM-Fit utilize VidePose3D for pose extraction, IMUTube incorporates extensive additional pre- and post-processing steps. This hypothesis is supported by the results for the video modality, where our trainable models outperform the baseline in all scenarios. Since the video embeddings, which we used as inputs, were extracted in the same manner for both our FiMAD dataset and MM-Fit, there is no such inconsistency, unlike the poses.
Leveraging multimodal inputs achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities. In this multimodal scenario, our method, compared to the baseline, improves the Macro $F_{1}$-Score from 0.824 to 0.882 (with NULL) and from 0.897 to 0.942 (without NULL) when trained on the full dataset. Furthermore, with only 2% of the training data available, we improve performance from 0.656 to 0.734 (with NULL) and from 0.790 to 0.885 (without NULL). These findings show that the pre-trained classifiers significantly outperform the baseline and learn more efficiently with a limited amount of training data across multiple input modalities.
{NiceTabular}
|c|c|ccc|ccc|
With “NULL” Input Modality Mean Macro $F_{1}$-Score
Train Data: $2\%$ Train Data: $100\%$
Baseline ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Baseline ([Uncaptioned image]) Pre-trained ([Uncaptioned image]) Pre-trained ([Uncaptioned image])
[…]
Table 2: Comparison of classification results for pose, video, and multimodal inputs evaluated on the MM-Fit test dataset when utilizing 2% and 100% of the training data. Each experiment was conducted 20 times.
Figure 4: Classification performance across various training data fractions on MM-Fit. The baseline model is compared to models with a pre-trained encoder and projection (both with frozen and trainable weights) on all input modalities (sensor, pose, video, and multimodal).
6.3 Zero-Shot
Table 6.3 demonstrates zero-shot results of MuJo on all four datasets. For MM-Fit and MotionSense, multimodal training outperforms unimodal training, achieving constantly higher top-1, top-3, and top-5 accuracy and Macro $F_{1}$-Scores. The MHEALTH dataset shows minimal differences between both training methods. However, for the MYOGYM dataset, which contains 30 gym exercises where all but 2 require equipment like machines, barbells, and more, mutimodal training performs worse. The reason for this might be that our pre-training was conducted on YouTube videos of fitness trainers primarily performing HIIT workouts, at home and not in the gym, focusing on bodyweight exercises. This domain shift may explain the decrease in performance from unimodal to multimodal training on MYOGYM: since the video features in our pre-training dataset lack representations of equipment-based exercises, they provide no valuable information for such types of exercises. Thus, the multimodal model might struggle more to recognize activities involving gym equipment. Consequently, unimodal training performs slightly better on this complex dataset, suggesting that target domain differences significantly impact performance.
The overall low scores are mainly due to data divergences between our pre-training data and the evaluation datasets, particularly for pose (MM-Fit) and sensor data. Since we pre-trained with virtual sensor data and poses extracted by IMUTube, our models struggle when applied to real sensor datasets and variations in pose representation (MM-Fit), leading to decreased performance. Video features perform the best out of all modalities because they were extracted in the same way during pre-training. Further, the models were not pre-trained with the NULL class, which is represented as the mean of all target class embeddings. Introducing this unknown class during evaluation increases classification complexity and leads to an additional drop in performance.
{NiceTabular}
|c|c|c|ccc|ccc|ccc|ccc| Dataset With “NULL” Input Modality Metrics
Accuracy   Macro $F_{1}$-Score
Unimodal Training   Multimodal Training   Unimodal Training   Multimodal Training
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
MM-Fit Sensor .112 .310 .491 .169 .378 .575 .084 .274 .461 .114 .341 .536
Pose .159 .369 .587 .196 .561 .671 .027 .233 .443 .061 .420 .574
Video .393 .565 .652 .501 .660 .733 .285 .494 .611 .427 .616 .712
Sensor .027 .396 .862 .064 .559 .879 .027 .165 .400 .036 .227 .498
Pose .035 .080 .127 .044 .282 .409 .006 .188 .370 .015 .301 .475
Video .073 .113 .454 .111 .477 .920 .107 .240 .395 .201 .417 .680
MHEALTH Sensor .117 .317 .515 .100 .335 .536 .093 .268 .454 .081 .273 .484
Sensor .161 .584 .817 .164 .638 .829 .054 .213 .418 .052 .203 .417
MYOGYM Sensor .083 .161 .235 .052 .122 .172 .054 .117 .190 .032 .093 .135
Sensor .021 .046 .147 .009 .026 .070 .021 .054 .113 .005 .041 .078
MOTIONSENSE Sensor .143 .428 .765 .146 .481 .867 .076 .353 .745 .104 .453 .850
Table 3: Consolidated zero-shot results contrasting uni- and multimodal training of MuJo on FiMAD.
7 Conclusion and Future Work
In this study, we introduced FiMAD, a multimodal dataset that contains over 78 hours of recordings. Our method, MuJo, is a novel pre-training approach that learns a joint feature space integrating video, pose, simulated accelerometer, and textual data. Our approach provides gains in real datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH, showing that our pre-training method captures a meaningful representation for HAR. Despite being trained solely on virtual IMU data, it enhances performance on real accelerometer data, especially when less data was used for training. Evenmore, our approach offers the flexibility to be applied not only to sensor inputs but also to video and pose inputs. Finally, although our zero-shot results are not outstanding, they demonstrate that multimodal training can enhance performance compared to unimodal training, if the target domain does not differ too much from our pre-training domain.
While our study presents a robust and innovative approach to HAR by integrating multiple data modalities into a joint feature space, there are several avenues for future research and enhancement. One potential improvement is to fine-tune the text and vision models during pre-training, rather than using pre-calculated text and video embeddings. Additionally, we plan to conduct a more detailed evaluation in the future, including fine-tuning MuJo on the target dataset using real sensor data, rather than relying solely on our pre-trained model. Furthermore, we aim to enhance the handling of the NULL class, which can be achieved by using a large language model to label the chapters of FiMAD that do not contain relevant activities as NULL. This would provide a clear representation for the NULL class in our dataset. Future work also includes incorporating other proxy tasks in our framework during pre-training as well as scaling the networks. In that setup we aim to train a single MuJo model instead of training a separate one for each target dataset when the sensor positions differ.